OpenStudy (anonymous):

5. By trial and improvement method,determine the roots of the following quadratic equations. @iambatman @wio @TheSmartOne @Jhannybean @perl

TheSmartOne (thesmartone):

waiting...

OpenStudy (anonymous):

a)x^2+5x+4=0

TheSmartOne (thesmartone):

umm.

TheSmartOne (thesmartone):

what 2 numbers when multiplied equal 4 and when you add tehm 5?

TheSmartOne (thesmartone):

them*

OpenStudy (anonymous):

\[2\times2=4\]

TheSmartOne (thesmartone):

try another pair.

TheSmartOne (thesmartone):

but 2x3=6 not 4

TheSmartOne (thesmartone):

there is one last pair...

TheSmartOne (thesmartone):

for example lets do it for 6 6=3x2 6=6x1

TheSmartOne (thesmartone):

that should give it away :P

OpenStudy (anonymous):

5-1=4 like this?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

wait what do you mean?

TheSmartOne (thesmartone):

what are two numbers that add up to 5 and multiple to 4. when there is a 1 difference you should realize that it is 1 and the number which is getting multiplied...

OpenStudy (anonymous):

1+4=5

TheSmartOne (thesmartone):

And 1x4=4

OpenStudy (anonymous):

thnx a lt @TheSmartOne

OpenStudy (anonymous):

lot*

TheSmartOne (thesmartone):

Well when we factor it we would have (x+1)(x+4)=0

TheSmartOne (thesmartone):

@MARC_ We are not done yet :P

TheSmartOne (thesmartone):

And then we set it up equal to 0 and we get x+1=0 x+4=0

TheSmartOne (thesmartone):

The fastest way would just be to graph it and see what the x intercept is. See: https://www.desmos.com/calculator/r2gwb3uljz

OpenStudy (anonymous):

x=-1 and x=-4

OpenStudy (anonymous):

thnx again @TheSmartOne

TheSmartOne (thesmartone):

No problem!

TheSmartOne (thesmartone):

And exactly!

OpenStudy (anonymous):

:)